# Code Graph Model (CGM): A Graph-Integrated Large Language Model for Repository-Level Software Engineering Tasks

### v2

@@ Abstract
Recent advances in Large Language Models (LLMs) have shown promise in function-level code generation, yet repository-level software engineering tasks remain challenging. Current solutions predominantly rely on proprietary LLM agents, which introduce unpredictability and limit accessibility, raising concerns about data privacy and model customization. This paper investigates whether open-source LLMs can effectively address repository-level tasks without requiring agent-based approaches. We demonstrate this is possible by enabling LLMs to comprehend functions and files within codebases through their semantic information and structural dependencies. To this end, we introduce Code Graph Models (CGMs), which integrate repository code graph structures into the LLM’s attention mechanism and map node attributes to the LLM’s input space using a specialized adapter. When combined with an agentless graph RAG framework, our approach achieves a 43.00% resolution rate on the SWE-bench Lite benchmark using the open-source Qwen2.5-72B model. This performance ranks first among open weight models, second among methods with open-source systems, and eighth overall, surpassing the previous best open-source model-based method by 12.33%. [^1] The code is available at https://github.com/codefuse-ai/CodeFuse-CGM.
**footnotetext: Equal Contribution.footnotetext: Correspondence to: Hang Yu <hyu.hugo@antgroup.com>, Jianguo Li <lijg.zero@antgroup.com>, Peng Di <dipeng.dp@antgroup.com>.
Keywords Code LLM  $\cdot$ Repository-Level  $\cdot$ Graph
@@ 1 Introduction
The dream of automating software engineering (SE) has long captivated both the SE and artificial intelligence (AI) communities [1, 2, 3]. Recent advancements in Large Language Models (LLMs) have shown promising results, particularly in code generation at the function level, with models achieving resolution rates above 90% on benchmarks such as HumanEval [4]. Unfortunately, real-world SE tasks extend far beyond isolated functions or self-contained code files. This is exemplified by repository-level issue resolution [5, 6], which encompasses not only software maintenance—addressing bugs and technical debt—but also software evolution, which involves introducing new features and enhancements [7].
@@
The complexity of repository-level coding tasks has led researchers and practitioners to assume that sophisticated strategies are necessary for their completion [8]. Indeed, current leading approaches typically utilize LLM agents powered by proprietary models like GPT-4/4o [9] and Claude 3.5 Sonnet [10]. These agents are designed to leverage tools, execute commands, observe environmental feedback, and plan subsequent actions [11]. Nevertheless, these methods suffer from two problems. First, the agent-driven mechanism introduces unpredictability in decision-making [2]. As the reasoning processes become intricate in tackling complex problems, the accumulation of errors can hinder the generation of optimal solutions [12]. Second, the reliance on closed-source models creates substantial barriers for the broader SE community [13, 14], including limited accessibility, inability to enhance or customize models for specific tasks, and serious security concerns regarding the privacy of sensitive code repositories when interacting with external API services.
@@
The above two challenges lead to a bold question: Can open-source LLMs be employed in an agentless manner to complete repository-level coding tasks? At first glance, this seems improbable. Closed-source agent-based approaches can resolve up to 55% of issues on the popular SWE-bench Lite benchmark [^2] https://www.swebench.com/ for issue fixing, whereas existing methods using open-source models have only achieved a maximum resolution rate of 30.67% as of May 2025 [15].
@@
Despite these initial reservations, we posit that the answer is “Yes”, and the key lies in empowering the open-source LLMs to fully comprehend code repositories, not just the information within individual functions and files, but also the dependencies across functions and files. To move forward to this goal, we propose Code Graph Models (CGMs), to jointly model the semantic and structural information of code repositories. Specifically, we first construct a code graph for each repository, which characterizes the hierarchical and reference dependencies between code entities. We then develop a method to integrate this graph into the LLM through two key mechanisms. (i) Semantic Integration: Node attributes (containing code or comments) are first encoded by a pretrained text encoder and then mapped to the LLM’s input space via an adapter, enabling the model to understand the semantic information of all nodes. (ii) Structural Integration: The graph structure is incorporated into the LLM through the attention mask, allowing direct message passing only between neighboring nodes in each layer of the LLM, similar to spatial Graph Neural Networks (GNNs) [16]. The entire system—comprising the text encoder, adapter, and LLM decoder—is then fine-tuned using Low Rank Adaptation (LoRA) [17]. The resulting CGM can tackle repository-level coding tasks by using both the code graph and user instructions (text format). To further augment the abilities of the CGM, we develop a specially designed Graph Retrieval-Augmented Generation (RAG) framework, consisting of four modules: Rewriter, Retriever, Reranker, and Reader (i.e., CGM). The first three modules focus the CGM on the subgraph that is most pertinent to the user’s query or issue.
@@
Our approach has demonstrated remarkable results on the SWE-bench Lite benchmark, reaching a 43.00% resolution rate using the open-source Qwen2.5-72B model and our agentless RAG framework. As of May 2025, this performance ranks first among methods utilizing open-source models, second among methods with open-source code implementations (the underlying model may still be closed-source), and eighth overall. Notably, our approach surpasses the previous best method based on open-source models (Moatless+DeepSeek-V3 [15]) by 12.33%, despite that method employing DeepSeek-V3, which shows stronger performance than Qwen2.5-72B.
@@ 2.1 Large Language Models for Code
Recent advancements in LLMs have shown remarkable success in generating code at self-contained function or file levels [3]. This includes powerful closed-source models like GPT-4/4o [9], Gemini-2.0 [18], and Claude 3.5 Sonnet [10], as well as open-source alternatives such as Llama 3.1 [19], Qwen 2.5 [20], and DeepSeek-V3 [21]. Additionally, code-specialized open-source models have also emerged, including StarCoder [22, 23], DeepSeek-Coder [14, 24], and Qwen-Coder [25]. However, these models struggle with repository-level coding tasks that better reflect practical software development scenarios. Even the most capable closed-source models achieve only modest success rates on the SWE-bench Lite benchmark [5] for real-world issue fixing, while open-source models lag further behind with a maximum resolution rate of 26% [26]. Although closed-source models show superior performance, their limited accessibility and data privacy concerns hinder widespread adoption in the SE community. Furthermore, their proprietary nature prevents fine-tuning on task-specific data to improve performance, if even such data is available.
For open-source LLMs to better handle repository-level tasks, they must develop a comprehensive understanding of both semantic and structural information within codebases. DeepSeek-Coder [14] has attempted to address this challenge by pre-training models on topologically sorted repository codes. However, this approach faces two major limitations: real-world repositories often contain more code than can fit within the model’s maximum context length; and the conversion of repository structure into text format tends to obscure explicit dependencies that exist in the codebase.
@@ 2.2 Graphs in Code Language Models
The integration of graph structures into code language models can be classified into three main approaches [27]: (1) attention mask modification, (2) graph-to-text conversion, and (3) positional encoding augmentation. In the first approach, models like GraphCodeBERT [28] and StructCoder [29] modify attention masks to capture relationships between code tokens in Abstract Syntax Trees (ASTs) and Data Flow Graphs (DFGs). The second approach, demonstrated by TreeBERT [30] and UniXcoder [31], transforms ASTs or node paths into textual sequences that can be processed by language models. The third approach, exemplified by TPTrans [32], leverages relative positional encodings to represent structural relationships within ASTs.
@@ 2.3 Agent-drive Methods for Software Engineering
LLM-based agents like Devin [33] have shown the potential to solve real-world SE problems through their reasoning [34, 35] and interactive capabilities [36, 37, 38, 11]. Along this direction, researchers have worked to enhance LLM agents through various approaches, including specialized agent-computer interfaces (ACI) [39, 40, 8, 41], fine-grained search [42, 12, 11], and expanded action spaces [43].
However, these agent-based approaches face several drawbacks. First, they typically delegate decision-making to the agents, allowing them to determine both the timing and nature of actions. While agents base their decisions on previous actions and environmental feedback, the expansive action space and complex feedback mechanisms can lead to repetitive behaviors or accumulating errors, ultimately resulting in suboptimal solutions [12]. Second, resolving a single issue often requires 30-40 interaction turns, making the process time-consuming and complicating the identification of specific turns that resulted in unsatisfactory outcomes [2]. Third, the inherent unpredictability of agent behavior and reliance on closed-source models creates obstacles for leveraging data to improve performance, despite the abundance of such data in practice, such as issue-patch pairs for issue fixing [5]. While SWE-Gym [44] attempts to address trainability, it may introduce bias by only training with the trajectories that lead the SWE agent to correct answers. As a remedy, we propose the CGM, built on open-source LLMs and enhanced through an agentless Graph RAG framework.
@@ 2.4 Agentless Methods for Software Engineering
Agentless models offer a more controlled approach to simulating real-world SE processes by following well-defined, fixed steps rather than relying on LLM agents to make autonomous decisions or use complex tools. They help avoid the issues of unpredictability and lengthy interaction chains. These methods typically operate in two main stages: localization and editing [45]. The localization stage identifies relevant code snippets within a repository, while the editing stage generates or modifies code based on these identified sections. This framework is particularly effective for repository-level code completion tasks, especially when combined with RAG [46, 47]. For more complex tasks like issue fixing, enhanced approaches with additional steps exist [2, 1]. For instance, Agentless [2] implements a comprehensive ten-step pipeline, dedicating four steps to improving localization accuracy. This method has achieved a promising resolution rate of 40.67% on SWE-bench Lite, comparable to state-of-the-art (SOTA) agent-based methods, though it relies on the closed-source model Claude-3.5 Sonnet.
Recent research has also focused on enhancing code understanding by incorporating structural information through graph-enhanced repository modeling [48, 49, 45]. However, even when graph structures are used during retrieval, existing methods typically flatten the retrieved code snippets into linear text sequences for downstream model prompting. This flattening process fails to preserve the inherent heterogeneity between graph and text modalities. As a remedy, we propose the CGM that explicitly aligns these two distinct modalities, enabling better preservation and utilization of structural information throughout the entire process.
@@ 3 Code Graph Construction
As shown in Figure 1, the hierarchical dependencies (i.e., the solid edges) span the code graph. In other words, all nodes are interconnected by edges reflecting hierarchical dependencies, establishing a top-down tree structure. This structure mirrors the organization of code entities as dictated by file systems and programming language syntax rules. Building this tree graph begins with AST parsing [48]. During this phase, code entities and their hierarchical dependencies are identified in a recursive manner: the root node (i.e., REPO) is added to the graph first, followed by its children (i.e., PKG and T-FILE), until all nodes without descendants (i.e., FUNC) are processed. With each recursion, directed edges are added from parents to children.
@@
Concerning node attributes, we retain the original content and line range of each node. This approach enables explicit graph traversal and retrieval and facilitates training models with enhanced semantic understanding capabilities. During post-processing, we remove the text contained in the child nodes from the parent nodes within the tree graph derived from the hierarchical dependencies. The resulting code graph is a text-rich graph [50] in which each node encapsulates a corresponding code snippet.
@@ 4.1 Model Architecture
Specifically for the encoder, we utilize the pretrained encoder from CodeT5+ [51], chosen for its proven effectiveness in processing both source code and text (comments and documentation). For nodes containing lengthy text, we segment the content into chunks of 512 tokens each. These chunks are processed independently by the encoder. To maintain graph consistency, we duplicate the source node for each chunk, preserving identical connections to other nodes. The chunks within a node are fully connected, and their sequential order is maintained through position embeddings in the LLM decoder ${\mathcal{D}}$. We fine-tune the encoder using Low-Rank Adaptation (LoRA) [17] to optimize its performance for downstream tasks.
The adapter $\mathcal{A}$ serves as a bridge between the encoder and LLM, projecting encoder outputs into the LLM’s input embedding space. Following successful practices in Vision Language Models (VLMs) [52, 53], we implement the adapter as a two-layer MLP with GELU activation [54]. The adapter is trained from scratch with random initialization.
Unlike VLMs, which bridge different modalities, CGM’s encoder ${\mathcal{E}}$ and decoder ${\mathcal{D}}$ are of the same modality, simplifying the alignment process. Furthermore, we compress each 512-token chunk (shown as gray tokens in Figure 2(d)) into a single node token (black tokens in Figure 2(d)) for the LLM decoder. This compression effectively extends the LLM’s context length by a factor of 512, enabling the processing of extensive code repository contexts. Similar techniques, referred to as soft prompt compression, have been shown to enhance long-context modeling in recent studies [55, 56, 57].
Structural Integration: Besides node information, another challenge is integrating the graph structure into the LLM decoder ${\mathcal{D}}$. While LLMs excel at processing sequential data, they are not inherently designed to capture graph structures [50]. Traditional approaches have attempted to incorporate repository-level structural information by simply linearizing code snippets into sequences [14, 45], but this transformation often fails to preserve the explicit relationships between code entities.
To better maintain structural relationships, we introduce a graph-aware attention mask to replace the causal attention mask solely between node tokens in the LLM. This mask is derived from the code graph’s adjacency matrix, taking into account the node duplication process described earlier. We then fine-tune the LLM with LoRA to adapt it to both the new attention mechanism and the node tokens from the adapter ${\mathcal{A}}$. This approach ensures that attention occurs only between neighboring nodes in the code graph, mimicking the message passing mechanism frequently used in spatial GNNs [58, 59].

### v3

@@ Abstract
Recent advances in Large Language Models (LLMs) have shown promise in function-level code generation, yet repository-level software engineering tasks remain challenging. Current solutions predominantly rely on proprietary LLM agents, which introduce unpredictability and limit accessibility, raising concerns about data privacy and model customization. This paper investigates whether open-source LLMs can effectively address repository-level tasks without requiring agent-based approaches. We demonstrate this is possible by enabling LLMs to comprehend functions and files within codebases through their semantic information and structural dependencies. To this end, we introduce Code Graph Models (CGMs), which integrate repository code graph structures into the LLM’s attention mechanism and map node attributes to the LLM’s input space using a specialized adapter. When combined with an agentless graph RAG framework, our approach achieves a 43.00% resolution rate on the SWE-bench Lite benchmark using the open-source Qwen2.5-72B model. This performance ranks first among open weight models, second among methods with open-source systems, and eighth overall, surpassing the previous best open-source model-based method by 12.33%.
Figure 1: Results on SWE-bench Lite. CGM-SWE-PY ranks first among open-weight models. CS-3.5 denotes Claude-3.5-Sonnet, DS-V3 represents DeepSeek-V3.
@@ 1 Introduction
The dream of automating software engineering (SE) has long captivated both the SE and artificial intelligence (AI) communities (Ma et al., 2024a; Xia et al., 2024; Zhang et al., 2023b). Recent advancements in Large Language Models (LLMs) have shown promising results, particularly in code generation at the function level, with models achieving resolution rates above 90% on benchmarks such as HumanEval (Chen et al., 2021). Unfortunately, real-world SE tasks extend far beyond isolated functions or self-contained code files. This is exemplified by repository-level issue resolution (Jimenez et al., 2024; Zan et al., 2024), which encompasses not only software maintenance—addressing bugs and technical debt—but also software evolution, which involves introducing new features and enhancements (Liu et al., 2024d).
@@
The complexity of repository-level coding tasks has led researchers and practitioners to assume that sophisticated strategies are necessary for their completion (Yang et al., 2024b). Indeed, current leading approaches typically utilize LLM agents powered by proprietary models like GPT-4/4o (OpenAI, 2024a) and Claude 3.5 Sonnet (Anthropic, 2024). These agents are designed to leverage tools, execute commands, observe environmental feedback, and plan subsequent actions (Ma et al., 2024b). Nevertheless, these methods suffer from two problems. First, the agent-driven mechanism introduces unpredictability in decision-making (Xia et al., 2024). As the reasoning processes become intricate in tackling complex problems, the accumulation of errors can hinder the generation of optimal solutions (Antoniades et al., 2024). Second, the reliance on closed-source models creates substantial barriers for the broader SE community (Manchanda et al., 2024; Zhu et al., 2024), including limited accessibility, inability to enhance or customize models for specific tasks, and serious security concerns regarding the privacy of sensitive code repositories when interacting with external API services.
@@
The above two challenges lead to a bold question: Can open-source LLMs be employed in an agentless manner to complete repository-level coding tasks? At first glance, this seems improbable. Closed-source agent-based approaches can resolve up to 55% of issues on the popular SWE-bench Lite benchmark [^1] https://www.swebench.com/ for issue fixing, whereas existing methods using open-source models have only achieved a maximum resolution rate of 30.67% as of May 2025 (Aor, 2024).
@@
Despite these initial reservations, we posit that the answer is “Yes”, and the key lies in empowering the open-source LLMs to fully comprehend code repositories, not just the information within individual functions and files, but also the dependencies across functions and files. To move forward to this goal, we propose Code Graph Models (CGMs), to jointly model the semantic and structural information of code repositories. Specifically, we first construct a code graph for each repository, which characterizes the hierarchical and reference dependencies between code entities. We then develop a method to integrate this graph into the LLM through two key mechanisms. (i) Semantic Integration: Node attributes (containing code or comments) are first encoded by a pretrained text encoder and then mapped to the LLM’s input space via an adapter, enabling the model to understand the semantic information of all nodes. (ii) Structural Integration: The graph structure is incorporated into the LLM through the attention mask, allowing direct message passing only between neighboring nodes in each layer of the LLM, similar to spatial Graph Neural Networks (GNNs) (Chami et al., 2022). The entire system—comprising the text encoder, adapter, and LLM decoder—is then fine-tuned using Low Rank Adaptation (LoRA) (Hu et al., 2021). The resulting CGM can tackle repository-level coding tasks by using both the code graph and user instructions (text format). To further augment the abilities of the CGM, we develop a specially designed Graph Retrieval-Augmented Generation (RAG) framework, consisting of four modules: Rewriter, Retriever, Reranker, and Reader (i.e., CGM). The first three modules focus the CGM on the subgraph that is most pertinent to the user’s query or issue.
@@
Our approach has demonstrated remarkable results on the SWE-bench Lite benchmark, reaching a 43.00% resolution rate using the open-source Qwen2.5-72B model and our agentless RAG framework. As of May 2025, this performance ranks first among methods utilizing open-source models, second among methods with open-source code implementations (the underlying model may still be closed-source), and eighth overall. Notably, our approach surpasses the previous best method based on open-source models (Moatless+DeepSeek-V3 (Aor, 2024)) by 12.33%, despite that method employing DeepSeek-V3, which shows stronger performance than Qwen2.5-72B.
@@ 2.1 Large Language Models for Code
Recent advancements in LLMs have shown remarkable success in generating code at self-contained function or file levels (Zhang et al., 2023b). This includes powerful closed-source models like GPT-4/4o (OpenAI, 2024a), Gemini-2.0 (Pichai et al., 2024), and Claude 3.5 Sonnet (Anthropic, 2024), as well as open-source alternatives such as Llama 3.1 (Dubey et al., 2024), Qwen 2.5 (Yang et al., 2024a), and DeepSeek-V3 (Liu et al., 2024a). Additionally, code-specialized open-source models have also emerged, including StarCoder (Li et al., 2023a; Lozhkov et al., 2024), DeepSeek-Coder (Zhu et al., 2024; Guo et al., 2024), and Qwen-Coder (Hui et al., 2024). However, these models struggle with repository-level coding tasks that better reflect practical software development scenarios. Even the most capable closed-source models achieve only modest success rates on the SWE-bench Lite benchmark (Jimenez et al., 2024) for real-world issue fixing, while open-source models lag further behind with a maximum resolution rate of 26% (Xie et al., 2025). Although closed-source models show superior performance, their limited accessibility and data privacy concerns hinder widespread adoption in the SE community. Furthermore, their proprietary nature prevents fine-tuning on task-specific data to improve performance, if even such data is available.
For open-source LLMs to better handle repository-level tasks, they must develop a comprehensive understanding of both semantic and structural information within codebases. DeepSeek-Coder (Zhu et al., 2024) has attempted to address this challenge by pre-training models on topologically sorted repository codes. However, this approach faces two major limitations: real-world repositories often contain more code than can fit within the model’s maximum context length; and the conversion of repository structure into text format tends to obscure explicit dependencies that exist in the codebase.
@@ 2.2 Graphs in Code Language Models
The integration of graph structures into code language models can be classified into three main approaches (Zhang et al., 2024c): (1) attention mask modification, (2) graph-to-text conversion, and (3) positional encoding augmentation. In the first approach, models like GraphCodeBERT (Guo et al., 2020) and StructCoder (Tipirneni et al., 2024) modify attention masks to capture relationships between code tokens in Abstract Syntax Trees (ASTs) and Data Flow Graphs (DFGs). The second approach, demonstrated by TreeBERT (Jiang et al., 2021) and UniXcoder (Guo et al., 2022), transforms ASTs or node paths into textual sequences that can be processed by language models. The third approach, exemplified by TPTrans (Peng et al., 2021), leverages relative positional encodings to represent structural relationships within ASTs.
@@ 2.3 Agent-drive Methods for Software Engineering
LLM-based agents like Devin (Cognition, 2023) have shown the potential to solve real-world SE problems through their reasoning (Luo et al., 2024; Wang et al., 2024b) and interactive capabilities (Hong et al., 2023; Kong et al., 2024; Xie et al., 2024; Ma et al., 2024b). Along this direction, researchers have worked to enhance LLM agents through various approaches, including specialized agent-computer interfaces (ACI) (Xue et al., 2023; Lee et al., 2024; Yang et al., 2024b; Zhang et al., 2024a), fine-grained search (Zhang et al., 2024b; Antoniades et al., 2024; Ma et al., 2024b), and expanded action spaces (Wang et al., 2024c).
However, these agent-based approaches face several drawbacks. First, they typically delegate decision-making to the agents, allowing them to determine both the timing and nature of actions. While agents base their decisions on previous actions and environmental feedback, the expansive action space and complex feedback mechanisms can lead to repetitive behaviors or accumulating errors, ultimately resulting in suboptimal solutions (Antoniades et al., 2024). Second, resolving a single issue often requires 30-40 interaction turns, making the process time-consuming and complicating the identification of specific turns that resulted in unsatisfactory outcomes (Xia et al., 2024). Third, the inherent unpredictability of agent behavior and reliance on closed-source models creates obstacles for leveraging data to improve performance, despite the abundance of such data in practice, such as issue-patch pairs for issue fixing (Jimenez et al., 2024). While SWE-Gym (Pan et al., 2024a) attempts to address trainability, it may introduce bias by only training with the trajectories that lead the SWE agent to correct answers. As a remedy, we propose the CGM, built on open-source LLMs and enhanced through an agentless Graph RAG framework.
@@ 2.4 Agentless Methods for Software Engineering
Agentless models offer a more controlled approach to simulating real-world SE processes by following well-defined, fixed steps rather than relying on LLM agents to make autonomous decisions or use complex tools. They help avoid the issues of unpredictability and lengthy interaction chains. These methods typically operate in two main stages: localization and editing (Ouyang et al., 2024). The localization stage identifies relevant code snippets within a repository, while the editing stage generates or modifies code based on these identified sections. This framework is particularly effective for repository-level code completion tasks, especially when combined with RAG (Shrivastava et al., 2023; Zhang et al., 2023a). For more complex tasks like issue fixing, enhanced approaches with additional steps exist (Xia et al., 2024; Ma et al., 2024a). For instance, Agentless (Xia et al., 2024) implements a comprehensive ten-step pipeline, dedicating four steps to improving localization accuracy. This method has achieved a promising resolution rate of 40.67% on SWE-bench Lite, comparable to state-of-the-art (SOTA) agent-based methods, though it relies on the closed-source model Claude-3.5 Sonnet.
Recent research has also focused on enhancing code understanding by incorporating structural information through graph-enhanced repository modeling (Ding et al., 2022; Liu et al., 2024c; Ouyang et al., 2024). However, even when graph structures are used during retrieval, existing methods typically flatten the retrieved code snippets into linear text sequences for downstream model prompting. This flattening process fails to preserve the inherent heterogeneity between graph and text modalities. As a remedy, we propose the CGM that explicitly aligns these two distinct modalities, enabling better preservation and utilization of structural information throughout the entire process.
@@ 3 Code Graph Construction
As shown in Figure 2, the hierarchical dependencies (i.e., the solid edges) span the code graph. In other words, all nodes are interconnected by edges reflecting hierarchical dependencies, establishing a top-down tree structure. This structure mirrors the organization of code entities as dictated by file systems and programming language syntax rules. Building this tree graph begins with AST parsing (Ding et al., 2022). During this phase, code entities and their hierarchical dependencies are identified in a recursive manner: the root node (i.e., REPO) is added to the graph first, followed by its children (i.e., PKG and T-FILE), until all nodes without descendants (i.e., FUNC) are processed. With each recursion, directed edges are added from parents to children.
@@
Concerning node attributes, we retain the original content and line range of each node. This approach enables explicit graph traversal and retrieval and facilitates training models with enhanced semantic understanding capabilities. During post-processing, we remove the text contained in the child nodes from the parent nodes within the tree graph derived from the hierarchical dependencies. The resulting code graph is a text-rich graph (Jin et al., 2024) in which each node encapsulates a corresponding code snippet.
@@ 4.1 Model Architecture
Specifically for the encoder, we utilize the pretrained encoder from CodeT5+ (Wang et al., 2023), chosen for its proven effectiveness in processing both source code and text (comments and documentation). For nodes containing lengthy text, we segment the content into chunks of 512 tokens each. These chunks are processed independently by the encoder. To maintain graph consistency, we duplicate the source node for each chunk, preserving identical connections to other nodes. The chunks within a node are fully connected, and their sequential order is maintained through position embeddings in the LLM decoder ${\mathcal{D}}$. We fine-tune the encoder using Low-Rank Adaptation (LoRA) (Hu et al., 2021) to optimize its performance for downstream tasks.
The adapter $\mathcal{A}$ serves as a bridge between the encoder and LLM, projecting encoder outputs into the LLM’s input embedding space. Following successful practices in Vision Language Models (VLMs) (Liu et al., 2024b; Wang et al., 2024a), we implement the adapter as a two-layer MLP with GELU activation (Hendrycks & Gimpel, 2016). The adapter is trained from scratch with random initialization.
Unlike VLMs, which bridge different modalities, CGM’s encoder ${\mathcal{E}}$ and decoder ${\mathcal{D}}$ are of the same modality, simplifying the alignment process. Furthermore, we compress each 512-token chunk (shown as gray tokens in Figure 3(d)) into a single node token (black tokens in Figure 3(d)) for the LLM decoder. This compression effectively extends the LLM’s context length by a factor of 512, enabling the processing of extensive code repository contexts. Similar techniques, referred to as soft prompt compression, have been shown to enhance long-context modeling in recent studies (Liao et al., 2024; Tan et al., 2024; Li et al., 2024).
Structural Integration: Besides node information, another challenge is integrating the graph structure into the LLM decoder ${\mathcal{D}}$. While LLMs excel at processing sequential data, they are not inherently designed to capture graph structures (Jin et al., 2024). Traditional approaches have attempted to incorporate repository-level structural information by simply linearizing code snippets into sequences (Zhu et al., 2024; Ouyang et al., 2024), but this transformation often fails to preserve the explicit relationships between code entities.
To better maintain structural relationships, we introduce a graph-aware attention mask to replace the causal attention mask solely between node tokens in the LLM. This mask is derived from the code graph’s adjacency matrix, taking into account the node duplication process described earlier. We then fine-tune the LLM with LoRA to adapt it to both the new attention mechanism and the node tokens from the adapter ${\mathcal{A}}$. This approach ensures that attention occurs only between neighboring nodes in the code graph, mimicking the message passing mechanism frequently used in spatial GNNs (Veličković et al., 2018; Shehzad et al., 2024).